My daughter has three boats, four cars, and seven planes in a box. If three are chosen at random and not replaced, what is the probability that the first two are planes and the last one is a boat?
Answer by stanbon(29603)

My daughter has three boats, four cars, and seven planes in a box. If three are chosen at random and not replaced, what is the probability that the first two are planes and the last one is a boat?
Prob of getting the two planes: [7C2]/[14C2]= 7*6/14*13=6/26=3/13
Prob of then getting the one boat: [3C1]/[12C1]=3/12=1/4
Prob(2 planes then 1 boat)= (3/13)*(1/4)= 3/52
Cheers,
Stan H.